The root mean square velocity of one mole of a monoatomic gas having molar mass M is Urms . What is the relation between average Kinetic Energy (E) of the gas and Urms ?

Grade:11

1 Answers

Arun
5 years ago
Dear Mohit
 
Urms = sqrt (3RT/M)
 
K.E. = 3/2 KT
 
Urms = sqrt (2E/M)
 
Hope it helps
 
Regards
Arun (askIITians forum expert)